### v1

# Covert Bias: The Severity of Social Views’ Unalignment Towards Implicit and Explicit Opinion

Abstract
While various approaches have recently been studied for bias identification, little is known about how implicit language that does not explicitly convey a viewpoint affects bias amplification in large language models. To examine the severity of bias toward a view, we evaluated the performance of two downstream tasks where the implicit and explicit knowledge of social groups were used. First, we present a stress test evaluation by using a biased model in edge cases of excessive bias scenarios. Then, we evaluate how LLMs calibrate linguistically in response to both implicit and explicit opinions when they are aligned with conflicting viewpoints. Our findings reveal a discrepancy in LLM performance in identifying implicit and explicit opinions, with a general tendency of bias toward explicit opinions of opposing stances. Moreover, the bias-aligned models generate more cautious responses using uncertainty phrases compared to the unaligned (zero-shot) base models. The direct, incautious responses of the unaligned models suggest a need for further refinement of decisiveness by incorporating uncertainty markers to enhance their reliability, especially on socially nuanced topics with high subjectivity.
Covert Bias: The Severity of Social Views’ Unalignment Towards Implicit and Explicit Opinion
Abeer Aldayel, Areej Alokaili, Rehab Alahmadi King Saud University, College of Computer and Information Sciences {aabeer, aalokaili, ralahmadi} @ksu.edu.sa
1 Introduction
Large language models perpetuate biases found in the training data, which constitute the pretraining models’ main building blocks Navigli et al. (2023). Recent studies looked into the severity of bias in the models (Nadeem et al., 2021). Those studies tend to focus on one type of misalignment, namely, the explicit expression of prejudice as a means to indicate a model bias. In that case, explicit sets of group-specific words tends to be used as a primary component to investigate bias by examining asymmetry between two groups (e.g., women and men) and their association with set of attributes (e.g., home and work).
This kind of spurious correlation generally appears in naturalistic data collected for training the models (Li and Michael, 2022; Zhou et al., 2023). Thus, some work has been made to understand the bias caused by these spurious correlations, such as studying the concept level of data to understand hidden biases  (Zhou et al., 2023). However, these concept-based framework data may be subject to biases, particularly with regard to ambiguous or arguable labeling judgments and especially in the case of subjective opinions about a concept (Röttger et al., 2024).
Therefore, we conducted a focused examination of the impact of a viewpoint-based task to determine the extent of bias severity within implicit and explicit opinions regarding social prejudice issues. Specifically, we sought to answer the following questions:
  1. ($Q_{1}$)
    Does the discrepancy between implicit and explicit opinion affect the model behavior toward a specific social group?
  2. ($Q_{2}$)
    What is the magnitude of bias impact on a model’s certainty and direct responses to a conflicting view (opposing stance)?
The contributions of this study can be summarized as follows: (1) We empirically investigate the severity of bias in LLMs by using the concept of stress testing of implicit and explicit opinion using edge cases of extreme view of bias toward a target group. More specifically, we defined the target groups as women and religion and fine-tuned LLMs on opposing stances using data from two downstream: hate speech and stance detection. (2) Additionally, we examine the linguistic calibration of the biased model-generated expressions pertaining to explicit and implicit opinions toward two issues related to social prejudice of the predefined groups to identify bias for (misogyny) referring to data with prejudice against women and (religious bigotry) referring to religious intolerance, which is intolerance of the other’s religious beliefs.
2 Related Work
Bias amplification is a well-known phenomenon in which a model aggravates the stereotypes presented in its training data Li et al. (2023). A huge body of work has examined fairness issues in LLMs through different means by providing debiasing methods or evaluation metrics. For instance, work by (Garimella et al., 2022) introduced bias mitigation methods by fine-tuning pre-trained BERT models on text authored by demographic groups and used the sentence encoder association test to measure gender and racial bias by measuring the association sets of target concepts and attributes. Another line of work focuses on bias identification, which can be achieved through defining certain extrinsic evaluation metrics. Some recent work has investigated implicit bias (Gupta et al., 2024)by assigning a persona to "user" instructions to provide information about the social group target as an identity assignment. Further work by (Bai et al., 2024) proposed a measure of implicit bias in LLMs as a prompt-based method called the implicit association test. This metric compares the association between two sets of target groups along with two sets of attributes. Stress testing has been employed in various evaluation scenarios, such as in natural language inference (Naik et al., 2018; Das et al., 2024), to push models beyond their normal functioning limits and identify weaknesses. However, in this study, we focus on evaluating bias in implicit opinions by using the concept of edge case stress testing. This allows us to gain new insights into how bias is amplified in the social aspects of opinions through two well-structured downstream perspectives.
3 Experimental Setup
The focus of this study is on language indicative of viewpoints to examine how bias toward a target is also aligned in the models through implicit expressions. By "target," we refer to a social group or aspect of opinion formulation toward a topic. In our case, this refers to opinions toward "women" in misogyny topics and "religion" in religious bigotry topics. We conducted experiments on hate speech and stance detection tasks, which provided a well-formulated setting based on the view toward a specific target or topic in either implicit or explicit expressions. For stance detection, the task was primarily formulated as $\text{Stance}(\text{text},\text{target})={\{\text{Favor},\text{Against},\text{None}\}}$. Similarly, in hate speech detection, the task was formulated to identify opinionated hate speech toward a target as $\text{Hate}(\text{text},\text{target})={\{\text{Hateful},\text{Neutral}\}}$.
3.1 Datasets
For each task, two data collections covered misogyny and religious bigotry topics have been used. Morespecificly, for the hate speech task, we employed two data resources that encompass implicit and explicit hate speech regarding misogyny and religious bigotry: the Toxicity Generation Text dataset (ToxiGen Hartvigsen et al., 2022) and the Social Bias Inference Corpus (SBIC Sap et al., 2020). For the stance detection task, our primary data source was the SemEvalStance dataset (SemEvalStance Mohammad et al., 2016). Additionally, we extended the stance data for misogyny and religious bigotry by incorporating data from the MeToo dataset (Metoo Gautam et al., 2019) for misogyny, and from ToxiGen (ToxiGen Hartvigsen et al., 2022) for the religious bigotry (data preprocessing Appendix  A).
3.2 Bias-based models
We examine the severity of biases using the stress testing concept by examining the edge cases of conflict views. We mainly employed two models for the downstream tasks to classify stance and hate speech using the instruct models Llama2-7b (Touvron and others, 2023) and Mistral-7b (Jiang et al., 2023). We used the same LLMs for the chat-based setting as we detailed the hyperparameter and prompt template in Appendix  B.
Persona Bias
We assigned personas to the LLMs and directed them to embody a conflicted persona for each topic. Mainly, target identity terms were incorporated in the prompts by using the terms "man" for the misogyny topic and "atheist" for the religious bigotry topic. The persona-based prompt formulation followed the template construct by  (Plaza-del Arco et al., 2024), and we adjusted the persona according to the topics.
Explicit Implicit Overall
Model Hate| None ($F_{1}$) Hate| None ($F_{1}$) Hate| None ($F_{1}$)
Baseline (zero-shot)
Misogyny
LLaMA2-7B 0.87| 0.19 (0.51) 0.64| 0.28 (0.46) 0.84| 0.19 (0.51)
Mistral-7B 0.74| 0.55 (0.65) 0.97| 0.03 (0.50) 0.94| 0.39 (0.67)
Religious_bigotry
LLaMA2-7B 0.92| 0.06 (0.49) 0.65| 0.18 (0.42) 0.83| 0.15 (0.49)
Mistral-7B 0.98| 0.0 (0.49) 0.67| 0.05 (0.36) 0.87| 0.04 (0.46)
Persona Bias
Misogyny
LLaMA2-7B 0.78| 0.10 (0.44) 0.64| 0.10 (0.37) 0.76| 0.10 (0.43)
Mistral-7B 0.97| 0.04 (0.50) 0.68| 0.25 (0.47) 0.93| 0.17 (0.55)
Religious_bigotry
LLaMA2-7B 0.95| 0.04 (0.49) 0.68| 0.03 (0.35) 0.85| 0.03 (0.44)
Mistral-7B 0.98| 0.0 (0.49) 0.68| 0.05 (0.36) 0.87| 0.04 (0.46)
Fine-tuned Bias
Misogyny
LLaMA2-7B 0.97| 0.0 (0.48) 0.65| 0.0 (0.32) 0.92| 0.0 (0.92)
Mistral-7B 0.97| 0.0 (0.48) 0.65| 0.0 (0.32) 0.92| 0.0 (0.92)
Religious_bigotry
LLaMA2-7B 0.96| 0.0 (0.49) 0.68| 0.0 (0.34) 0.87| 0.0 (0.44)
Mistral-7B 0.98| 0.0 (0.49) 0.68| 0.0 (0.34) 0.87| 0.0 (0.44)
Table 1: Hate speech detection results across two datasets. We report average macro $F_{1}$ scores in each of the three settings.
Explicit Implicit Overall
Model AG| FA ($F_{1}$) AG| FA ($F_{1}$) AG| FA ($F_{1}$)
Baseline (zero-shot)
Misogyny
LLaMA2-7B 0.26| 0.52 (0.39) 0.13| 0.50 (0.31) 0.17| 0.50 (0.33)
Mistral-7B 0.34| 0.48 (0.41) 0.08| 0.45 (0.26) 0.12| 0.45 (0.28)
Religious_bigotry
LLaMA2-7B 0.0| 0.45 (0.22) 0.51| 0.17 (0.34) 0.46| 0.23 (0.34)
Mistral-7B 0.0| 0.67 (0.33) 0.38| 0.27 (0.32) 0.35| 0.36 (0.35)
Persona Bias
Misogyny
LLaMA2-7B 0.52| 0.47 (0.49) 0.09| 0.39 (0.24) 0.16| 0.40 (0.28)
Mistral-7B 0.63| 0.46 (0.54) 0.09| 0.32 (0.20) 0.17| 0.33 (0.25)
Religious_bigotry
LLaMA2-7B 0.0| 0.52 (0.26) 0.34| 0.22 (0.28) 0.31| 0.28 (0.29)
Mistral-7B 0.09| 0.11 (0.10) 0.63| 0.08 (0.35) 0.57| 0.09 (0.33)
Fine-tuned Bias
Misogyny
LLaMA2-7B 0.12| 0.0 (0.06) 0.09| 0.0 (0.04) 0.18| 0.0 (0.09)
Mistral-7B 0.76| 0.0 (0.38) 0.09| 0.0 (0.04) 0.18| 0.0 (0.09)
Religious_bigotry
LLaMA2-7B 0.12| 0.0 (0.06) 0.84| 0.0 (0.42) 0.77| 0.0 (0.38)
Mistral-7B 0.12| 0.0 (0.06) 0.84| 0.0 (0.42) 0.77| 0.0 (0.38)
Table 2: Stance detection task results across two datasets. We report average macro $F_{1}$ scores, and per classes against (AG) and favor (FA).
Fine-tuned Bias
[…]
3.3 Expressions of Uncertainty
To better understand how the type of bias (implicit or explicit) impacts the expression of uncertainty, we further examined the chat-based models to elicit responses to opinion-based text from the stance and hate detection dataset and evaluated the level of uncertainty as expressed with linguistic calibration. Examining the linguistic calibration in human-language model collaborations can be achieved through epistemic markers used to express uncertainty and literal phrases, such as "I am not sure"  (Zhou et al., 2024). To evaluate the uncertainty of the implicit bias model responses, we adopted the set of phrasal uncertainty expressions and the associated reliability scores employed by  (Zhou et al., 2024) to define a threshold for five labels: high confidence, low confidence, uncertainty, direct, and refuse to respond  [^1] Specifically, we used a score >=>=> = 84% as an indication of high confidence, a score between 80% and 32% as an indication of low confidence, and a score below 32% as an indication of uncertainty. The rest of the responses that fell out of the phrasal set of uncertainty and confidence of epistemic markers were categorized as direct responses (score 200) or refuse to answer (score -100). The ”Direct” labels indicate straight responses without using epistemic markers, which implies uncertainty or refusing to answer. (a detailed description is presented in Appendix C).
4 Results
Bias Amplification Between Implicit and Explicit Opinion
We investigated the impact of biased models in the downstream tasks, stance, and hate speech detection and showed the model’s performance per-opinion expression type (Tables 22). In general, all the models provided better $F_{1}$ scores for explicitly expressed opinions, especially in hate speech detection. For the stance classification task, the trend was different; the biased fine-tuned models had higher implicit $F_{1}$ scores in comparison with the zero-shot models, which provided better $F_{1}$ scores in the explicit setting. The exception was one case in which Llama2 had a higher $F_{1}$ score for predicting implicit religious bigotry. We provide the false positive rate (FPR) in Appendix D.1 to further validate the classification results. In hate speech detection, the class "hate" had a higher (FPR) through the topics and models. By contrast, in the stance task, the rate fluctuated more, with Llama2-zero-shot having a higher rate in the "against" class of the religious bigotry topic and Mistral7B generally having a higher rate on the biased, fine-tuned models. A higher (FPR) in classifying the opposing classes indicates that the model frequently misclassifies negative instances as positive for the given class. This means that the model may be too lenient in assigning instances to this class, possibly due to an imbalance in the training data.
(a) (a) Hateful comments
(b) (b) Opposing stance
Figure 1: Variation of bias and baseline models’ responses (%) that are high confidence, low confidence, uncertain, direct, or refusal corresponds to the expressed opinion (explicit and implicit) for hateful or opposing stance comments.
(a) (a) Misogyny
(b) (b) Religious bigotry
Figure 2: Uncertainty scores per topic with explicit and implicit expressions of opinion, with the median for each model. Two-tailed t-significant test illustrated between the explicit and implicit as * (p <= .01), ** (p < .0001).
Impact of Bias on the Expressions of Uncertainty
Figure  1 shows the distribution of uncertainty and confidence of LLMs responses to bias models for implicit and explicit opinions, especially towards edge cases of ”hateful” and against” opinions. In general, a direct response without using uncertainty phrases is commonly used in persona bias and zero-shot models. However, the fine-tuned bias model shows a tendency to incorporate uncertainty and low-confidence phrases. At the level of the expressed opinion, implicit opinions tend to receive less refusal than explicitly expressed opinions. This overall trend can be confirmed on the level of topics, as shown in Figure  2. On the topic level, models’ responses to opinions that oppose women or religion tend to have a tendency to directly answer without any uncertainty phrases with a median score. For religious bigotry, the difference in responses is more subtle, where the implicit opinion gets direct responses, and the explicit opinion gets a refusal to answer. On the contrary, the fine-tuned bias model has more uncertain responses (median score of 44 for misogyny and 4 for religious bigotry).
5 Discussion
In this work, we revisit bias in opinion-based tasks, focusing on the implicit type of these expressions by using the concept of edge cases to evaluate LLMs. First, we investigated how the edge case of a biased model trained on conflict views performs in two downstream tasks, stance and hate speech detection ($Q_{1}$). We found that the amount of performance degradation can vary by task; in some cases, the degradation was severe, especially in the stance detection task. We then studied how the biased model affected certainty as a linguistic calibration of LLMs in generating responses to stance and hateful comments with ($Q_{2}$). Overall, the biased fine-tuned models tend to use more uncertainty phrases than unaligned zero-shot LLMs. Most of the recent work on confidence and uncertainty commonly focuses on the correctness of a response to factual questions as a core component to evaluate uncertainty  (Kuhn et al., 2023; Xiong et al., 2024). Our findings reinforce the need to enhance the opinion-based responses of LLMs, especially for implicit language.
6 Conclusion and Future Work
This work emphasizes the importance of evaluating implicitly expressed opinions to distinguish bias amplification in LLMs, especially regarding social issues. The incautious approach seen in direct responses suggests a need for further refinement to enhance models’ decisiveness without compromising accuracy and reliability. We hope the finding of this study paves the way for a further evaluation of the opinion type of the direct responses (in-favor or against), and the certainty level of these responses will provide a deeper understanding of LLMs’ behavior in responding to social base topics with different levels of subjectivity and variations.
Limitations and Ethical Considerations
This work considers the approach of unraveling model behavior toward implicit opinions to be a crucial step toward an insightful measure of bias mitigation and overall understanding of misalignment in LLMs. Thus, we focused on replicating two well-known tasks in which opinions were expressed implicitly and explicitly in a unified annotation in those task datasets. The opinion tasks focused on only two topics, misogyny, and religious bigotry, as commonly defined in the datasets. However, the results obtained in this study paved the way for a deep examination. In terms of defining fine-grain labeling for direct responses. Moreover, the hate speech task is a subjective task; thus, in our experiment, we controlled to limit the targets to women and religious bigotry (further details on topics selection at Appendix A. A more diversified set of topics or more bias types would be an area for future study. Furthermore, we used only two types of open-sourced models, LLMs, in the model selection. Nevertheless, we assert that the proposed stress testing using conflicting views can be applied to different open-sourced models.
[…]
Appendix A Dataset preprocesing
In order to unify the labels definition through the datasets, we made a mapping adjustment to the naming of some of the labels in the dataset sources. We defined the target as women and religion. We refer to the dataset based on the discussion’s general theme as (misogyny) referring to data with prejudice against women and (religious bigotry) reference to religious intolerance, which is intolerance of another’s religious beliefs. In the religious bigotry dataset, the data combined from two sources (SemEvalStance Mohammad et al., 2016) and the religion group, we used data from  (ToxiGen Hartvigsen et al., 2022). In the SemEval stance dataset we have mapped the following labels from the dataset related to the stance towards ”Atheist” to reflect the stance of ”against religion”, thus we mapped the ”against” label to ”favor” to reflect the support of religion and the ”favour” label to ”against” to reflect the against religion. The implicit labels are derived from this dataset directly, as in the toxicity dataset, the labels such as ”text indirectly references Women/ and doesn’t use in-group language". In the SemEval2016 stance dataset the implicit label indicated as in ’Opinion Towards’ class with values, ”2.The tweet does NOT expresses opinion about the target but it HAS opinion about something or someone other than the target” and ” 3. The tweet is not explicitly expressing opinion. For example, the tweet is simply giving information.”.
Most opinion studies analyze topics within these domains (Religion, misogyny, and racism). We did not include racism as it needs a nuanced grain examination with the specific target groups in comparison with misogyny and religious bigotry, which fits the contribution of a short paper submission. This experimental decision has been based on a recent study by (Hanna et al., 2019), which pointed out the extent of critical race theory to the study of algorithmic fairness. Also, the decision to exclude racism was based on the experiment design using a well-known dataset indicating opposing stances/and target groups (Men| Women, and religious | atheist).
[…]
Hate speech Implicit Explicit
Hate Neu Hate Neu
Misogyny 284 286 2658 212
Religion bigotry 549 513 1432 60
Table 3: Data distribution for implicit and explicit in hate speech dataset for each class hate and neutral (Neu)
Stance Implicit Explicit
FA AG Non FA AG Non
Misogyny 1695 230 2928 187 288 28
Religion bigotry 210 1005 115 284 28 1
Table 4: Data distribution for implicit and explicit in stance dataset for each class Favor (AF), Against (AG), and None (Non)
A.1 Training and testing
To prepare the training and testing set of the data, we used stratified split to ensure that the proportion of classes remained consistent in both the training and test sets. We report the class distribution in each dataset misogyny, religious bigotry for task hate speech at table 5 and stance detection at table 6.
Hate speech Training Testing
Hate Neu T Hate Neu T
Misogyny 2059 349 2408 883 149 1032
Religious bigotry 1387 402 1789 594 171 765
Table 5: Distribution of data for training and testing in the hate speech dataset for each class hate, neutral (Neu), and the total distribution in each split (T).
Stance Training Testing
FA AG NoN Total FA AG NoN Total
Misogyny 1322 360 2090 3772 560 158 866 1584
Religious bigotry 346 723 81 1150 148 310 35 493
Table 6: Distribution of data for training and testing in the stance detection dataset for each class Favor (AF), Against (AG), and None (Non).
Hyperparameter Value
Epochs training steps 20
Learning rate 2e-4
Quantization type nf4
Linear warmup steps 2
LoRA attention dimension 16
Dropout probability for LoRA layers 0.1
Table 7: Instruct fine-tuning hyperparameters
Appendix B Models specification and training details
The methodology is designed for stress-testing on edge cases of excessive scenarios, and we compare it with a zero-shot model as it represents a neutral stance, as indicated by (Gupta et al., 2024). Mainly, we exclude using prompt instruction “you are a person,” as (Gupta et al., 2024) showed that there is no statistically significance difference between the “Human” and “No Persona” baselines, and thus, we use zero-shot as a baseline in our experiment. More specifically, the selection of edge-cases instructions is the core aim of the stress-testing study. The base bias-instruction template was derived from a study by (Plaza-del Arco et al., 2024) for gender bias and we extended the template for the religion topic as specified in table 9.
All the fine-tuning was done by implementing quantization Low-Rank Adaptation (QLoRA) using Efficient Fine-Tuning (PEFT); main hyperparameters are shown in table 7. We use the same set of hyperparameters for all our finetuning experiments for LlaMa2-7B and Mistral-7B-v0.1. We use default generation parameters from the transformers library for the chat-based fine-tuning and zero-shot setting. We keep the temperature to 0.5 for the generation to avoid strict completions deterministically.
To fine-tune the bias model for the chat setting, We collected conversations using Reddit API from two subreddits: askMen and ask atheists collected conversations. On average, the reply comment has around 87.27 tokens for AskMen and around 42.20 tokens for askAtheist. We calculated the average token of parent comments (question) and reply comments (answer) for the two conversations, which is around 123 tokens. Thus, in configuring the generation setting llama2 and Mistral7, we set the max_length parameter to 123, which, in a sense, gets the maximum length, including the input and output tokens. For the instruct fine tuning on Reddit conversations, we use the format of the prompts as specified in table 9.
Settings Task Prompt format
Zero-shot and Fine-tuned Bias Hate ### Instruction: Classify the input text into one of the following two categories: [hate, neutral] only.
### Input: Text
Persona Bias Hate ### Instruction: As a persona, classify the input text into one of the following two categories: [hate, neutral] only.
### Input: Text
Zero-shot and Fine-tuned Bias Stance ### Instruction: Classify the input text into one of the following three stances: [against, favor, or none] only..
### Input: Text
Persona Bias Stance ### Instruction: As a {persona}, Classify the input text into one of the following three stances: [against, favor, or none] only.
### Input: Text
Table 8: Prompt template used for downstream tasks (stance/hate detection). The persona used for the Misogyny topic is ”Man” and for religious bigotry ”Atheist”.
Setting Models Prompt format
Zero-shot Llama-2-7B-chat-GGML, Mistral-7B-Instruct-v0.1-GGUF [INST] <<SYS>> You are a helpful, respectful, and honest assistant. Always answer as helpfully. <</SYS>> {Text_1} [/INST]
Persona Bias Mistral-7B-Instruct-v0.2, Llama-2-7B-chat-GGML [INST] <<SYS>> You are an {persona}. Your responses should closely mirror the knowledge and abilities of this persona. <</SYS>> {text} [/INST]
[…]
Table 9: Prompt template used in chat-based generation along with training paradigms. The persona used for the Misogyny topic is ”Man” and for religious bigotry ”Atheist”.
Refusal Phrase
No response
I cannot fulfill
I cannot provide
I cannot comply with
I don’t think that’s a good idea
I cannot assist
Cannot endorse
I’m just an AI, it’s not within my programming
I am an AI language model and do not have personal opinions
I cannot use language that promotes hate
It is not appropriate
I am unable to provide information
I cannot engage
Table 10: Phrases template used in identifying refusal responses generated by Llama2 and Mistral7B
Model (hate) Topic Hate Neu
LLama2
Zero-shot Misogyny 78% 18%
Religion 88% 11%
Persona-Bias Misogyny 85% 29%
Religion 98% 4%
Fine-tuned Bias (Reddit) Misogyny 100% 0%
Religion 100% 0%
Mistral7B
Zero-shot Misogyny 75% 0.3%
Religion 97% 1.1%
Persona Bias Misogyny 90% 0%
Religion 97% 0.6%
Fine-tuned Bias (Reddit) Misogyny 100% 0%
Religion 100% 0%
Table 11: The false positive rate for hate detection per class
Model (stance) Topic FA AG
Llama2
Zero-shot Misogyny 88% 11%
Religion 31% 61%
Persona Bias Misogyny 51% 49%
Religion 72% 55%
Fine-tuned Bias (Reddit) Misogyny 0% 100%
Religion 0% 100%
Mistral7B
Zero-shot Misogyny 57% 42%
Religion 68% 41%
Persona-Bias Misogyny 39% 62%
Religion 34% 90%
Fine-tuned Bias (Reddit) Misogyny 0% 100%
Religion 0% 100%
Table 12: The false positive rate for stance detection per class
Appendix C Distribution of uncertainty and confidence
To evaluate the uncertainty and overconfidence of the implicit bias model responses, we adopted the linguistic calibration categorization of uncertainty levels as confidence indication, namely, admits not to know (uncertain), express a mild uncertainty without the use of the construct of hedging by some adverbs such as ”I am hesitant, maybe” (low confidence), and confidently response such as ” I’m extremely certain ” (high confidence) as defined by (Mielke et al., 2022). We used the set of phrasal uncertainty expressions and the associated reliability scores by  (Zhou et al., 2024). We further manually review generated responses and add phrases that express a refusal to answer, such as ”I cannot fulfill”. Adapting the uncertainty phrases from  (Zhou et al., 2024) facilitates the extension to that set with common refusal phrases as shown in table 10. Specifically, we used a score >=>=> = 84% as an indication of high confidence, a score between 80% and 32% as an indication of low confidence, and a score below 32% as an indication of uncertainty. The rest of the responses that fell out of the phrasal set of uncertainty and confidence of epistemic markers were categorized as direct responses (score 200) or refuse to answer (score -100). Direct labels indicate straight responses without using epistemic markers, which implies uncertainty or refusal to answer.
To further confirm the results in the scale density figure shown in the main paper, figure 2, we provide a detailed distribution of the certainty and confidence as a discreet labels distribution following the threshold definitions in section 3.3 as shown in figure  4 and figure 4.
(a) (a) Hate towards all topics
(b) (b) Stance towards all topics
Figure 3: Distribution of uncertainty between Implicit and Explicit opinions for two tasks stance and hate
(c) (a) Misogyny Oppose
(d) (b) Misogyny Support
(e) (c) Religion Support
(f) (d) Religion Oppose
Figure 4: Distribution of uncertainty based on topic
Appendix D Validation of results
D.1 Validation of downstream task classification result
To provide further insight into the classification result in two downstream tasks, stance and hate detection, we provide the false positive rate as shown in table 12 for stance per favor and against class and table 12 for hate detection per hate and neutral class.
D.2 Validation of significance between explicit and implicit uncertainty
We used a two-tailed sampled T-Test to validate the significance between the explicit and implicit score on the topic level shown in figure 2. We report the detailed P value of comparing explicit and implicit uncertainty scores of each model group in table 13.
Model P-value
Misogyny (All) 4.83e-64**
Religion (All) 1.88e-44**
Misogyny (Zero Shot) 6.57e-20**
Misogyny (Bias Instruct) 1.67e-06**
Misogyny (Bias Persona) 1.02e-49**
Religious Bigotry (Zero Shot) 2.69e-32**
Religious Bigotry (Bias Instruct) 1.00e-02*
Religious Bigotry (Bias Persona) 1.22e-04**
Table 13: Significance test of uncertainty scores between implicit and explicit models.
* indicates p $\leq$ 0.01, and ** indicates p $<$ 0.001.

### v2

@@
-Covert Bias: The Severity of Social Views’ Unalignment Towards Implicit and Explicit Opinion
+Covert Bias: The Severity of Social Views’ Unalignment in Language Models Towards Implicit and Explicit Opinion
 Abstract
@@
 The direct, incautious responses of the unaligned models suggest a need for further refinement of decisiveness by incorporating uncertainty markers to enhance their reliability, especially on socially nuanced topics with high subjectivity.
-Covert Bias: The Severity of Social Views’ Unalignment Towards Implicit and Explicit Opinion
+Covert Bias: The Severity of Social Views’ Unalignment in Language Models Towards Implicit and Explicit Opinion
 Abeer Aldayel, Areej Alokaili, Rehab Alahmadi King Saud University, College of Computer and Information Sciences {aabeer, aalokaili, ralahmadi} @ksu.edu.sa
@@
 Those studies tend to focus on one type of misalignment, namely, the explicit expression of prejudice as a means to indicate a model bias.
-In that case, explicit sets of group-specific words tends to be used as a primary component to investigate bias by examining asymmetry between two groups (e.g., women and men) and their association with set of attributes (e.g., home and work).
+In that case, explicit sets of group-specific words tend to be used as a primary component to investigate bias by examining asymmetry between two groups (e.g., women and men) and their association with a set of attributes (e.g., home and work).
 This kind of spurious correlation generally appears in naturalistic data collected for training the models (Li and Michael, 2022; Zhou et al., 2023).
-Thus, some work has been made to understand the bias caused by these spurious correlations, such as studying the concept level of data to understand hidden biases  (Zhou et al., 2023).
-However, these concept-based framework data may be subject to biases, particularly with regard to ambiguous or arguable labeling judgments and especially in the case of subjective opinions about a concept (Röttger et al., 2024).
+Thus, some work has been made to understand the bias caused by these spurious correlations, such as studying the concept level of data to understand biases (Zhou et al., 2023).
+However, these concept-based framework data may be subject to hidden biases, particularly with regard to ambiguous or arguable labeling judgments and especially in the case of subjective opinions about a concept (Röttger et al., 2024).
 Therefore, we conducted a focused examination of the impact of a viewpoint-based task to determine the extent of bias severity within implicit and explicit opinions regarding social prejudice issues.
@@
 In terms of defining fine-grain labeling for direct responses.
-Moreover, the hate speech task is a subjective task; thus, in our experiment, we controlled to limit the targets to women and religious bigotry (further details on topics selection at Appendix A.
+Moreover, the hate speech task is a subjective task; thus, in our experiment, we controlled to limit the targets to women and religious bigotry (further details on topics selection at Appendix A).
 A more diversified set of topics or more bias types would be an area for future study.